# From Lab to Field: Real-World Evaluation of an AI-Driven Smart Video Solution to Enhance Community Safety

### v2

@@ Introduction
Overall, these real-world evaluations allows research how these key parameters, such as the number of cameras and crowd density, could affect system performance, such as latency and throughput in different public safety scenarios in case of emergency, is crucial in decision making and taking the proper actions. For example, retail and transport hubs could analyze foot traffic and crowd density, offering insights into peak hours and emergency response planning (Hosseini et al. 2024).
Conclusion
@@
Our evaluation highlighted the system’s ability to maintain operational efficiency with increasing node counts, despite a corresponding rise in PCP latency. Object and behavioral anomaly detection latencies ranged from 4.7 seconds with four cameras to 26.76 seconds with sixteen, reflecting the system’s scalability and robustness. Through extensive testing, the system demonstrated its capability to deliver timely notifications to end users, crucial for managing anomalies in public spaces. This work underscores the importance of real-world evaluations in optimizing AI-driven video solutions, ensuring they meet the demands of dynamic environments and contribute to enhanced public safety and operational efficiency.

### v3

@@ 7.3 Physical-Cyber-Physical Evaluation (Anomaly detection)
Overall, these real-world evaluations allow research into how key parameters—such as the number of cameras and crowd density—affect system performance metrics like latency and throughput. Understanding these dynamics in different public safety scenarios, especially in cases of emergency, is crucial for informed decision-making and timely action. For instance, in retail environments and transportation hubs, the system can provide valuable insights into foot traffic patterns and crowd density fluctuations, facilitating both operational optimization and emergency response. [39]
8 Conclusion
@@
Our evaluation highlighted the system’s ability to maintain operational efficiency with increasing node counts, despite a corresponding rise in PCP latency. Object and behavioral anomaly detection latencies ranged from 4.7 seconds with four cameras to 26.76 seconds with sixteen, reflecting the system’s scalability and robustness. Through extensive testing, the system demonstrated its capability to deliver timely notifications to end users, which is crucial for managing anomalies in public spaces. This work underscores the importance of real-world evaluations in optimizing AI-driven video solutions, ensuring they meet the demands of dynamic environments and contribute to enhanced public safety and operational efficiency.